A card is chosen at random from a deck of 52 cards. Its is not replaced and a second card is chosen.What is the probability of getting a jack then an eight?
Answer by sudhanshu_kmr(1152)


probability of getting jack on first chance = 4/52
probability of getting an eight on second chance = 4/51

probability = 4/52 *4/51
= 4/ 663